Ammonia is produced synthetically by the reaction below. How many moles of NH3 are formed when 200.0 g of N2 reacts with hydrogen?
1
Expert's answer
2013-01-16T10:22:23-0500
3H2 + N2 = 2NH3
First, are looking for the amount of substanceof nitrogen:
n(N2) = 200 g / 28 g/mol = 7.14 mol
n(NH3) = n(N2) * 2
n(N2) = 7б14 * 2 = 14.28 mol
Answer:
n(NH3) = 14.28 mol